# The Energy Cost of Artificial Intelligence of Things Lifecycle

### v2

@@ IX Conclusions and Future Work
In this work, we proposed a novel metric, namely, eCAL. Unlike the traditional metrics that focus only on the energy required for transmitting, compute infrastructure or for AI models, eCAL can capture the overall energy cost of generating an inference in an AIoT system during the entire lifecycle of a trained model. We propose a detailed methodology to determine the eCAL of an AIoT system by breaking it down into various data manipulation components, such as data collection, preprocessing, training, evaluation, and inference, and analyzing the complexity and energy consumption of each component. The proposed metric demonstrates that the more a model is utilized, the more energy-efficient each inference becomes. For example, considering a simple MLP architecture, the energy consumption per bit for 100 inferences is $2.73$ times higher than for 1000 inferences. Furthermore, we developed an open-source modular energy estimation tool based on our approach, enabling fine-grained analysis of energy consumption across AIoT components, allowing developers to evaluate and optimize AIoT energy efficiency in various deployment scenarios.
Through the proposed eCAL metric, this study provides foundations for understanding the energy consumption of AIoT system architectures. Future work may focus on extending the study from AIoT system to other communication system architectures that are increasingly relying on AI techniques for their automation and performance optimization, especially the emerging open radio access network (O-RAN) based cellular systems. The exploration of the relationship between model performance and eCAL for 6G verticals and related network slices may also be worthy of pursuing. Finally, the development of dynamic energy management algorithms in view of a more sustainable operation is highly relevant.

### v3

@@ IX Conclusions and Future Work
In this work, we have proposed a novel metric, namely, eCAL. Unlike traditional metrics that focus only on energy required for transmission, computing infrastructure, or AI models separately, eCAL can capture the overall energy cost of generating inferences in communication system during the entire lifecycle of a trained model. We have proposed a detailed methodology to determine the eCAL of an AI enhanced communication system by breaking it down into various data manipulation components, such as data collection, preprocessing, training, evaluation, and inference, and analyzing the complexity and energy consumption of each component. The proposed metric demonstrates that the more a model is utilized, the more energy-efficient each inference becomes. For example, considering a simple MLP architecture, the energy consumption per bit for 100 inferences is $2.73$ times higher than for 1000 inferences. Furthermore, we have developed an open-source, modular energy estimation tool based on our approach. This tool enables fine-grained analysis of energy consumption across components, allowing developers to evaluate and optimize AI energy efficiency in various deployment scenarios.
Through the proposed eCAL metric, this study lays the foundation for understanding the energy consumption of AI-enhanced emerging communication architectures. Future work may focus on extending this study beyond the assumptions, case studies, and configurations provided in this study to cover existing and emerging communication system designs that are increasingly relying on AI techniques for their automation and performance optimization, especially the emerging open radio access network (O-RAN) based cellular systems. The exploration of the relationship between model performance and eCAL for 6G verticals, related network slices, online, distributed, fine tuned and federated model lifecycles may also be worthy of pursuing. Finally, the development of dynamic energy management algorithms in view of a more sustainable operation is highly relevant.